OpenStudy (syahirahshmsl):

how to solve sin ( theta + pi/2 ) = 1/2 ?

OpenStudy (danjs):

you remember the angle sum identity, sin(x + y)=sin(x)cos(y) + cos(x)sin(y)

OpenStudy (danjs):

Using that for the sin(theta + pi/2) it will be the same as cos(theta) so you get \[\cos(\theta)=1/2\]

OpenStudy (danjs):

there are two angles where the cosine is 1/2...